please help! Chapter 6, Section 2-CI, Exercise 086 Use the t-distribution to find a confidence interval for a mean y given the relevant sa
Answer #1

Point estimate=\bar{X}=85.9

Now , df=degrees of freedom=n-1=15-1=14

The critical value is , tdf,a/2 = t14,0.01 /2 = 2.977 ; From t-table

The margin of error is ,

E=t_{df,\alpha/2}\frac{s}{\sqrt{n}}=2.977*\frac{30.0}{\sqrt{15}}=23.06

The 99% confidence interval is ,

\bar{X}\pm E

85.9\pm 23.06

(62.84,108.96)